A metal rod has a length of 100cm at 200°C. At what temperature will its lenght be 99.4cm. If the linear expansivity of the material of the rod is 150.0002/K


1
Expert's answer
2021-01-31T18:57:44-0500
"\dfrac{\Delta L}{L_0}=\alpha (T-T_0),""T=\dfrac{L-L_0}{\alpha L_0}+T_0,""T=\dfrac{99.4\ cm-100\ cm}{2.0\cdot10^{-5}\ \dfrac{1}{^{\circ}C}\cdot100\ cm}+200\ ^{\circ}C=-100^{\circ}C."